# Towards Open Vocabulary Learning: A Survey

### v1

@@ 3.3 Open Vocabulary Object Detection
In this section, we review the approaches for open vocabulary object detection. We divide the methods into six different categories: knowledge distillation, region text pertaining, training with more balanced data, prompting modeling, and region text alignment.
@@
Knowledge Distillation. The aim of these techniques is to distill the knowledge of Vision-and-Language Models (VLMs) into close-set detectors [121, 2, 4]. Since the knowledge of VLMs is much larger than close-set detectors, the distillation of novel classes into based classes trained detectors is a straightforward idea. One earlier method is ViLD [13], a two-stage detection approach that utilizes instance-level visual-to-visual knowledge distillation. ViLD consists of two branches: the ViLD-text branch and the ViLD-image branch. In the former branch, instead of using a learnable classifier, fixed text embeddings are employed, which are obtained from the output of VLMs’ text encoder. Meanwhile, in the latter branch, pre-computed proposals are first fed into a detector to obtain region embeddings. The cropped image is then sent to VLMs’ image encoder to generate image embeddings. Subsequently, ViLD proposes distilling this information onto each Region-of-Interest (RoI) via $\mathcal{L}_{1}$ Loss. HierKD [157] uses a single-stage detector and introduces a global-level language-to-visual knowledge distillation module that aims to narrow down performance gaps between one-stage methods and two-stage methods. This technique employs global-level knowledge distillation modules (GKD), which align global-level image representations with caption embeddings through contrastive loss. Both ViLD and HierKD use pixel-based detectors. However, Rasheed et al. [181] leverages query-based detector Deformable DETR [148] for its detection process. To ensure consistency between their detection region representations and CLIP’s region representations, the authors employed inter-embedding relationship matching loss (IRM). Furthermore, the authors also adopt mixed datasets pre-training to enhance the ability of novel class discovery. OADP [182] thinks previous works only distill object-level information from VLMs to downstream detectors and ignore the relation between different objects. To tackle this problem, OADP employs object-level distillation as well as global and block distillation methods. These supplementary techniques aim to compensate for the lack of relational information in object distillation by optimizing the L1 distance between CLIP visual encoder and the detector backbone’s global features or block features. Rather than using simple novel class names for text distillation, there are also several works utilizing more fine-grained information, including attributes, captions, and relationships of objects for distillation. PCL [183] adopts an image captioning model to generate captions that describe object instances in a more comprehensive way. OVRNet [184] simultaneously detects objects and their visual attributes in open vocabulary scenarios. By exploring COCO attributes dataset [185] via a joint co-training strategy, the authors find that the recognition of fine-grained attributes works complementary for OVD.
@@
Region Text Pertaining. Another assumption of open vocabulary learning is the availability of large-scale image text pairs, which can be easily obtained in daily life. Since these pairs contain large enough knowledge, which can cover the most novel or unseen datasets for detection and segmentation. The learning of region text alignment maps the novel classes of visual features and text features into an aligned feature space. OVR-CNN [12] first introduces the concept of open vocabulary object detection by using caption data for novel class detection. The model first trains a ResNet[1] and vision to language (V2L) layer using image-caption pairs via grounding, masked language modeling, and image-text matching. Since captions are not constrained in language space, the V2L layer learns to map features from visual space into semantic space without limiting to closed-label space. During the next stage of training, Faster R-CNN [121] is used as a detection method with pre-trained ResNet as its backbone. OVR-CNN [12] replaces only learnable classifiers with fixed text embeddings from pre-trained language models [9]. For classification purposes, region visual features obtained from RoI-Align [2] are sent into the V2L layer and mapped into semantic space. Attribute-Sensitive OVR-CNN [186] proposes a different approach from OVR-CNN [12]. Instead of grounding vision regions to the input word embeddings of BERT [9], Attribute-Sensitive OVR-CNN aligns vision regions with contextualized word embeddings that are output from BERT [9]. Additionally, Attribute-Sensitive OVR-CNN proposes using an adjective-noun negative caption sampling strategy to enhance the model’s sensitivity to words such as adjectives, verb phrases, and prepositional phrases other than object nouns in the caption. RegionCLIP [159] learns visual region representation by matching image regions to region-level descriptions. It creates pseudo labels by CLIP for region-text pairs and then uses contrastive loss to match them before fine-tuning the visual encoder using human-annotated detection datasets. OWL-ViT [187] removes the final token pooling layer of pre-trained VLMs’ image encoder and attaches a lightweight classification head and box regression head to each transformer output token before fine-tuning it on standard detection datasets using bipartite matching loss. Meanwhile, MaMMUT [168] presents a simple text decoder and visual encoder for multimodal pertaining. By designing a two-pass on text decoder, it combines both contrastive and generative learning in one framework. The former for grounding text visual entities while the latter learns to generate.
@@
Training with More Balanced Data. Rare and unseen data are common in image classification datasets. Joint training can be used to address this issue. The core idea of these approaches by leveraging more balanced data, including image classification datasets, pseudo labels from image-text data, or extra-related detection data. Detic [160] improves long-tail detection performance with image-level supervision. The classification head of Detic is trained using image-level data from ImageNet21K [188]. During training, the max area proposal from RPN is chosen as the RoI. The basic insight of Detic is that most classification data are object-centric. Therefore, the maximum area proposal may completely cover an object that’s represented by an image-level class. The mm-ovod [189] method improves upon Detic [160] by utilizing multi-modal text embeddings as the classifier. This approach employs a large language model to create a description of each class name, rather than relying on manual prompt templates for generating text-based embeddings. In addition to this, mm-ovod also generates vision-based embeddings from image exemplars. By fusing the text-based and vision-based embeddings, the multi-modal text embeddings can significantly enhance Detic’s performance. To use more data, several methods generate pseudo bounding box annotations from large-scale image-caption pairs. PB-OVD [190] uses Grad-CAM [191] and generates the pseudo-bounding boxes in combination with the proposals from RPN. The activation map of Grad-CAM is obtained from the alignment between region embeddings and word embeddings that come from pre-trained VLMs. Then the boxes are generated from such map and jointly trained with existing box annotations. Meanwhile, several works leverage the rich semantics available in recent vision and language models to localize and classify objects in unlabeled images. VL-PLM [158] proposes to train Faster R-CNN as a two-stage class-agnostic proposal generator using a detection dataset without category information. LocOV [192] uses class-agnostic proposals in RPN to train Faster R-CNN by matching the region features and word embeddings from image and caption, respectively. From the data generation view, several works [193, 194, 195] adopt the diffusion model to generate the on-target data for effective training. In particular, X-Paste [193] generates the rare class data to improve the classification ability for existing approaches. It uses an extra segmentation model to provide the foreground object masks and adopt simple copy and paste [196] to augment training data. Recently, several works [167, 197, 198] have explored the self-training approaches to generate large pseudo labels for more balanced learning. OWLv2 [167] presents a self-training pipeline. It generates huge pseudo-box annotations on WebLI [199] dataset and pre-trains a model on such generated dataset. Finally, it fine-tunes this model on a specific OVD dataset. Since the WebLI contains a large vocabulary size, OWLv2 achieves significant gains on both LVIS and COCO datasets.
@@
Prompting Modeling. Prompt modeling is an effective technique for adapting foundation models to various domains, such as language modeling [200] and image classification [201, 202]. By incorporating learned prompts into the foundation model, the model can transfer its knowledge to downstream tasks more easily. To generate text embeddings of category names, prompts are fed to the text encoder of pre-trained VLMs. However, negative proposals do not belong to any specific category. To address this issue, DetPro [161] forces the negative proposal to be equally dissimilar to any object class instead of using a background class. PromptDet [36] introduces category descriptions into the prompt and explores the position of the category in the prompt. It also proposes to use cached web data to enhance the novel classes during training. Based on the DETR [4] framework, CORA [163] proposes region prompting and anchor pre-matching. The former reduces the gap between the whole image and region distributions by prompting the region features of the CLIP-based region classifier, while the latter learns generalizable object localization via a class-aware matching mechanism. Prompt-OVD [164] follows the pipeline of OV-DETR [162]. It presents RoI-based masked attention and RoI pruning techniques by utilizing CLIP visual features to improve the novel object classification.
Region Text Alignment. Using language as supervision instead ground truth bounding box is an attractive alternative for open vocabulary object detection. However, obtaining enough object-language annotations is difficult and costly. OV-DETR[162] introduces a transformer-based detector for open vocabulary object detection by replacing the bipartite matching method with a conditional binary matching mechanism. VLDet[35] converts the image into a set of regions and the caption into a set of words and solves the object-language alignment problem using a set matching method. It proposes a simple matching strategy to align caption and vision features. DetCLIPv2 [203] uses ATSS [204] as an object detector and trains it with three datasets: a standard detection dataset, a grounding dataset, and an image-text pairs dataset for word-text alignment.
@@
BARON [166] proposes to align the embedding in bags of different regions rather than only individual regions. It first groups contextually interrelated regions as a bag and treats each region in the bag as a word in a sentence. Then, it sends the bag of regions into the text encoder to get bag-of-regions embeddings. These bag-of-regions embeddings will align with cropped region embeddings from the image encoder of VLMs. F-VLM [165] finds that the origin CLIP features already have grouping effects. It is a two-branch method similar to ViLD-text [13]. F-VLM uses a CLIP vision encoder as the backbone and applies the VLM feature pooler on the region features from the backbone to get VLM predictions. The final result of F-VLM combines the detection scores and the VLM predictions. RO-ViT [205] builds upon F-VLM [165], which believes that the difference in position embeddings between image-level and region-level is responsible for the gap between vision-language pre-training and open vocabulary object detection. To address this issue, RO-ViT proposes a cropped positional embedding module in VLM that bridges the gap between vision-language pre-training and downstream open vocabulary object detection tasks.
@@ 3.4 Open Vocabulary Segmentation
In this section, we review the approaches for open vocabulary segmentation problems. Although segmentation problems can be defined in various ways, such as semantic segmentation or panoptic segmentation, we categorize the methods based on their technical aspects.
@@
Utilizing VLMs to Leverage Recognition Capabilities. VLMs have shown remarkable performance in image classification by learning rich visual and linguistic representations. Therefore, it is natural to extend VLMs to semantic segmentation, which can be seen as a dense classification task. For instance, LSeg [15] aligns the text embeddings of category labels from a VLM language encoder with the dense embeddings of the input image. This enables LSeg to leverage the generalization ability of VLMs and segment objects that are not predefined but depend on the input texts. Following LSeg, several works propose methods to utilize the VLMs for open vocabulary segmentation tasks. Fusioner [206] utilizes self-attention operations to combine visual and language features in a transformer-based framework at an early stage. ZegFormer [169] decouples the problem into a class-agnostic segmentation task and a mask classification task. It uses the label embeddings from a VLM to classify the proposal masks and applies the CLIP-vision encoder to obtain language-aligned visual features for them. After that, more approaches are proposed to extract the rich knowledge in VLMs. SAN [180] attaches a lightweight side network to the pre-trained VLM to predict mask proposals and classification outputs. CAT-Seg [179] jointly aggregates the image and text embeddings of CLIP by fine-tuning the image encoder. Han et al. [207] develop an efficient framework that does not rely on the extra computational burden of the CLIP model. OPSNet [208] proposes several modulation modules to enhance the information exchange between the segmentation model and the VLMs. The modulation modules fuse the knowledge of the learned vision model and VLM, which improves the zero-shot performance in novel classes. MaskCLIP [39] inserts Relative Mask Attention (RMA) modules into a pre-trained CLIP model. It can utilize the CLIP features more efficiently. TagCLIP [209] proposes a trusty token module to explicitly predict pixels that contain objects (both base and novel) before classifying each pixel, which avoids the problem that models tend to misidentify pixels into novel categories.
@@
Learning from Caption Data. Image caption data can provide weak supervision for identifying novel classes, as they expose potential novel category names. Like such approaches in open vocabulary object detection [12], image caption data is also explored in open vocabulary segmentation tasks. OpenSeg [37] applies a region-word grounding loss between mask features and caption word features, which enables it to achieve good zero-shot testing ability on different datasets. CGG [38] combines caption grounding and caption generation losses to fully exploit the knowledge in caption data. It leverages the role of object nouns in visual grounding and the mutual benefits of words in caption generation.
@@
Generating Pseudo Labels. Intuitively, providing the model with more data on novel categories can improve classification performance. MaskCLIP+ [170] modifies the CLIP-vision model by replacing the last pooling layer with a convolution layer, which produces dense feature maps. These feature maps are then used to generate pseudo labels for training a segmentation model. OVSeg [171] matches the proposed image regions with nouns in captions using CLIP to generate pseudo annotations. It also proposes a mask prompt tuning module to help CLIP adapt to masked images without changing their weights. XPM [210] follows a similar pseudo data generation procedure by aligning words in captions with regions in images to generate pseudo instance mask labels.
@@
Training without Pixel-Level Annotations. Despite the absence of annotations on novel objects, most methods still require base mask annotations during training, which are costly and labor-intensive to obtain. To reduce the annotation burden, many recent works explore training segmentation models with merely weak supervision, such as image caption. GroupViT [172] proposes a semantic segmentation framework that leverages a grouping mechanism to automatically merge image patches with the same semantics. It trains the model with a contrastive loss on image-caption pairs and does not need pixel-level annotations. PACL [211] enhances the contrastive loss with a patch alignment objective that aligns the image patches and the CLS token of the captions. ViL-Seg [173] combines both contrastive loss and clustering loss. SegCLIP [212] further introduces a reconstruction loss and a superpixel-based KL loss. To compute the superpixel-based KL loss, it uses an unsupervised graph-based segmentation model [213] to generate pseudo labels. TCL [174] proposes a finer-grained contrastive loss, namely text-grounded contrastive loss, which explicitly aligns captions and regions. The model can segment the region that corresponds to a given text expression during inference. OVSegmentor [178] introduces masked entity completion and cross-image mask constituency tasks to improve the training efficiency. Mask-free OVIS [214] generates pseudo mask annotations with purely image-text pair data to assist the segmentation models.
@@
Jointly Learning Several Tasks. Open vocabulary segmentation encompasses different tasks, including OVSS, OVIS, and OVPS. How to jointly learn multiple segmentation tasks in one model becomes a practical problem. X-Decoder [215] proposes a framework that can handle various tasks, including open vocabulary semantic segmentation, open vocabulary instance segmentation, and open vocabulary panoptic segmentation. It utilized a query-based segmentation architecture and pre-trains the model on a mixture of segmentation data and image-text pairs and then fine-tuned or applied in zero-shot settings for downstream tasks. FreeSeg [216] also proposes a generic framework to tackle the three tasks in a unified manner. It designs an adaptive task prompt module and performs test time tuning on the learnable prompts to capture the task-specific features. PROP [217] first trains class prompts at a large vocabulary dataset (Imagenet-21K) and then transfer the prompts into multiple open vocabulary tasks. Moreover, in case the model can learn from multiple tasks, it raises the question of whether the model can benefit from multi-sourced data, e.g., detection and segmentation data. To handle this question, OpenSeeD [177] jointly learns from segmentation and detection datasets. To bridge the task gap, they propose a decoupled decoding framework, which decodes foreground and background masks separately and generates masks for the bounding box proposals.
@@
Adopting Denoising Diffusion Models. Recently, diffusion-based generative models [218] have achieved remarkable success in text-based image generation suggests. There are mainly two ways to let diffusion models enhance open vocabulary tasks. First, The reality and diversity of the images generated by diffusion models suggest that the intermediate representations in the diffusion models may be highly aligned with natural language vocabularies. Inspired by this, ODISE [175] proposes a framework that leverages the middle representation of the diffusion model. The middle representations contain rich semantic information for both base and novel classes. Thus ODISE can perform open vocabulary segmentation by training a decoder head on the representations. Another way to incorporate diffusion models into the open vocabulary setting is to take advantage of the image generation ability. OVDiff [176] proposes a prototype-based method to tackle open vocabulary semantic segmentation. It uses the diffusion model to generate images for various categories and treat them as prototypes. The method does not need training. During testing, the input images are compared with these generated prototypes, and the best-matching one is the predicted class.

### v4

@@ 3.3 Open Vocabulary Object Detection
In this section, we divide the methods into five categories: knowledge distillation, region text pre-training, training with more balanced data, prompting modeling, and region text alignment. Finally, we summarize the common features in Tab. II.
@@
Knowledge Distillation. These techniques aim to distill the knowledge of Vision-and-Language Models (VLMs) into close-set detectors [157, 2, 4]. Since the knowledge of VLMs is much larger than close-set detectors, distilling novel classes into based classes trained detectors is a straightforward idea. Knowledge distillation aims to distill visual knowledge directly into close-set detectors since the visual features are aligned with text features during the VLM pre-training stage. One earlier method is ViLD [12], a two-stage detection approach that utilizes instance-level visual-to-visual knowledge distillation. ViLD consists of two branches: the ViLD-text branch and the ViLD-image branch. In the former branch, fixed text embeddings obtained from VLMs’ text encoder output are treated as classifiers. Meanwhile, in the latter branch, pre-computed proposals are fed into a detector to obtain region embeddings using the RoIAlign [2] function. The cropped image is then sent to the VLMs’ image encoder to generate image embeddings. Subsequently, ViLD proposes distilling this information onto each Region-of-Interest (RoI) via $\mathcal{L}_{1}$ Loss. LP-OVOD [223] extends the ViLD framework by making two main modifications. Firstly, it replaces the softmax cross entropy loss with the sigmoid focal loss [159]. Secondly, LP-OVOD introduces a new classification branch that is supervised by pseudo labels. HierKD [224] uses a single-stage detector and introduces a global-level language-to-visual knowledge distillation module. The module aims to narrow down performance gaps between one-stage and two-stage methods. This technique employs global-level knowledge distillation modules (GKD), which align global-level image representations with caption embeddings through contrastive loss. Both ViLD and HierKD use pixel-based detectors. However, Rasheed et al. [225] leverages query-based detector Deformable DETR [185] for its detection process. To ensure consistency between their detection region representations and CLIP’s region representations, the authors employed inter-embedding relationship matching loss (IRM). Furthermore, the authors adopt mixed datasets pre-training to enhance the ability of novel class discovery. OADP [226] thinks previous works only distill object-level information from VLMs to downstream detectors and ignore the relation between different objects. To tackle this problem, OADP employs object-level distillation as well as global and block distillation methods. These supplementary techniques aim to compensate for the lack of relational information in object distillation by optimizing the L1 distance between the CLIP visual encoder and the detector backbone’s global features or block features. Rather than using simple novel class names for text distillation, several works also utilize more fine-grained information, including attributes, captions, and relationships of objects. PCL [227] adopts an image captioning model to generate more comprehensive captions that describe object instances. OVRNet [228] simultaneously detects objects and their visual attributes in open vocabulary scenarios. By exploring the COCO attributes dataset [229] via a joint co-training strategy, the authors find that the recognition of fine-grained attributes works complementary for OVD. In summary, knowledge distillation is a common design, which effectively transfer VLM’s knowledge into close set detectors. However, the recognition ability is still within the scope of teacher VLMs.
@@
Region Text Pre-training. Another assumption of open vocabulary learning is the availability of large-scale image text pairs, which can be easily obtained in daily life. Since these pairs contain large enough knowledge to cover the most novel or unseen datasets for detection and segmentation. Most approaches adopt web-scale caption data for pre-training, which contains millions of image text pairs. The learning of region text alignment maps the novel classes of visual features and text features into an aligned feature space. Once trained for the alignment, it is nature to generalize the detector for novel class classification. OVR-CNN [11] first introduces the concept of open vocabulary object detection by using caption data for novel class detection. The model first trains a ResNet[1] and vision to language (V2L) layer using image-caption pairs via grounding, masked language modeling, and image-text matching. Since captions are not constrained in language space, the V2L layer learns to map features from visual space into semantic space without limiting to closed-label space. During the next stage of training, Faster R-CNN [157] is used as a detection method with pre-trained ResNet as its backbone. OVR-CNN [11] replaces only learnable classifiers with fixed text embeddings from pre-trained language models [9]. For classification purposes, region visual features obtained from RoI-Align [2] are sent into the V2L layer and mapped into semantic space. Attribute-Sensitive OVR-CNN [230] proposes a different approach from OVR-CNN [11]. Instead of grounding vision regions to the input word embeddings of BERT [9], Attribute-Sensitive OVR-CNN aligns vision regions with contextualized word embeddings that are output from BERT [9]. Additionally, Attribute-Sensitive OVR-CNN proposes using an adjective-noun negative caption sampling strategy to enhance the model’s sensitivity to adjectives, verb phrases, and prepositional phrases other than object nouns in the caption. GLIP series [231, 232] unify the object detection and phrase grounding for pre-training. In particular, it leverages massive image-text pairs by generating grounding boxes in a self-training fashion, which sets strong results for both detection and grounding. RegionCLIP [216] learns visual region representation by matching image regions to region-level descriptions. It creates pseudo labels by CLIP for region-text pairs and then uses contrastive loss to match them before fine-tuning the visual encoder using human-annotated detection datasets. OWL-ViT [233] removes the final token pooling layer of pre-trained VLMs’ image encoder and attaches a lightweight classification head and box regression head to each transformer output token before fine-tuning it on standard detection datasets using bipartite matching loss. Meanwhile, MaMMUT [220] presents a simple text decoder and visual encoder for multimodal pre-training. Designing a two-pass text decoder combines both contrastive and generative learning in one framework. The former is for grounding text visual entities, while the latter learns to generate. DITO [234] presents a new image-level pretraining strategy to bridge the gap between image-level pretraining and open-vocabulary object detection. At the pertaining phase, DITO replaces the classification architecture used in CLIP with the detector architecture, which better serves the region-level recognition needs of detection by enabling the detector heads to learn from noisy image-text pairs. In summary, adopting more text-image pair can improve the performance on rare and novel classes. However, such process needs more computation cost for extra dataset training.
@@
Training with More Balanced Data. Rare and unseen data are common in image classification datasets. Joint training can be used to address this issue. The core idea of these approaches is to leverage more balanced data, including image classification datasets, pseudo labels from image-text data, extra-related detection data, or even data generated by generation models. Detic [134] improves long-tail detection performance with image-level supervision. The classification head of Detic is trained using image-level data from ImageNet21K [235]. During training, the max area proposal from RPN is chosen as the RoI. Detic’s basic insight is that most classification data are object-centric. Therefore, the maximum area proposal may completely cover an object represented by an image-level class. The mm-ovod [236] method improves upon Detic [134] by utilizing multi-modal text embeddings as the classifier. This approach employs a large language model to create a description of each class to generate text-based embeddings. In addition, mm-ovod also uses vision-based embeddings from image exemplars. By fusing the text-based and vision-based embeddings, the multi-modal text embeddings can significantly enhance Detic’s performance. To use more data, several methods generate pseudo bounding box annotations from large-scale image-caption pairs. PB-OVD [119] uses Grad-CAM [237] and generates the pseudo-bounding boxes combined with RPN’s proposals. The activation map of Grad-CAM is obtained from the alignment between region embeddings and word embeddings that come from pre-trained VLMs. Then, the boxes are generated from the activation map and jointly trained with existing box annotations. Meanwhile, several works leverage the rich semantics available in recent vision and language models to localize and classify objects in unlabeled images. VL-PLM [215] proposes to train Faster R-CNN as a two-stage class-agnostic proposal generator using a detection dataset without category information. LocOV [238] uses class-agnostic proposals in RPN to train Faster R-CNN by matching the region features and word embeddings from image and caption, respectively. From the data generation view, several works [239, 240, 138] adopt the diffusion model to generate the on-target data for effective training. In particular, X-Paste [239] generates the rare class data to improve the classification ability for existing approaches. It uses an extra segmentation model to provide the foreground object masks and adopt simple copy and paste [241] to augment training data. Recently, several works [135, 242, 136] have explored the self-training approaches to generate large pseudo labels for more balanced learning. OWLv2 [135] presents a self-training pipeline. It generates huge pseudo-box annotations on WebLI [243] dataset and pre-trains a model on such generated dataset. Finally, it fine-tunes this model on a specific OVD dataset. Since the WebLI contains a large vocabulary size, OWLv2 achieves significant gains on LVIS and COCO datasets. In summary, these approaches are more effective in unbalanced data setting. Designing a more efficient data augmentation still have room to explore.
@@
Prompting Modeling. Prompt modeling is an effective technique for adapting foundation models to various domains, such as language modeling [244] and image classification [245, 246]. By incorporating learned prompts into the foundation model, the model can transfer its knowledge to downstream tasks more easily. To generate text embeddings of category names, prompts are fed to the text encoder of pre-trained VLMs. However, negative proposals do not belong to any specific category. To address this issue, DetPro [217] forces the negative proposal to be equally dissimilar to any object class instead of using a background class. PromptDet [45] introduces category descriptions into the prompt and explores the position of the category in the prompt. It also proposes to use cached web data to enhance the novel classes during training. Based on the DETR [4] framework, CORA [219] proposes region prompting and anchor pre-matching. The former reduces the gap between the whole image and region distributions by prompting the region features of the CLIP-based region classifier, while the latter learns generalizable object localization via a class-aware matching mechanism. Prompt-OVD [247] follows the pipeline of OV-DETR [218]. It presents RoI-based masked attention and RoI pruning techniques by utilizing CLIP visual features to improve the novel object classification. Despite the effectiveness, without more data training, the performance of prompting is limited compared with other directions.
@@
Region Text Alignment. Using language as supervision instead of a ground truth bounding box is an attractive alternative for open vocabulary object detection. However, obtaining enough object-language annotations is difficult and costly. Compared with region text pre-training, region text alignment aims at a better matching between region visual features and text features during the base class training without introducing extra data. OV-DETR[218] introduces a transformer-based detector for open vocabulary object detection by replacing the bipartite matching method with a conditional binary matching mechanism. VLDet[44] converts the image into a set of regions and the caption into a set of words, and solves the object-language alignment problem using a set matching method. It proposes a simple matching strategy to align caption and vision features. DetCLIPv2 [248] uses ATSS [249] as an object detector and trains it with three datasets: a standard detection dataset, a grounding dataset, and an image-text pairs dataset for word-text alignment. Recently, BARON [113] proposes to align the embedding in bags of different regions rather than only individual regions. It first groups contextually interrelated regions as a bag and treats each region in the bag as a word in a sentence. Then, it sends the bag of regions into the text encoder to get bag-of-regions embeddings. These bag-of-regions embeddings will align with cropped region embeddings from the image encoder of VLMs. CoDet [250] reformulates the region-word alignment as a co-occurring object discovery problem and aligns the co-occurring objects with the shared concept. F-VLM [137] finds that the origin CLIP features already have grouping effects. It is a two-branch method similar to ViLD-text [12]. F-VLM uses a CLIP vision encoder as the backbone and applies the VLM feature pooler on the region features from the backbone to get VLM predictions. The final result of F-VLM combines the detection scores and the VLM predictions. RO-ViT [251] builds upon F-VLM [137], which believes that the difference in position embeddings between image-level and region-level is responsible for the gap between vision-language pre-training and open vocabulary object detection. To address this issue, RO-ViT proposes a cropped positional embedding module in VLM that bridges the gap between vision-language pre-training and downstream open vocabulary object detection tasks. In summary, region-text alignment is a core research topic, and it still has room to explore when considering recent large language models [43].
TABLE III: Feature summarization of current open vocabulary segmentation approaches.
Method Common Feature
Utilizing VLMs to Leverage Recognition Capabilities. Design fusion or alignment methods to better fuse VLM knowledge into the existing segmenters.
Learning from Caption Data. Use extra caption data to ground the objects’ visual features with caption data.
Generating Pseudo Labels. Fully explore the potential of the VLM model to generate better masks to train the segmentation models
Training without Pixel-Level Annotations. Combine the VLM and different unsupervised approaches to perform unsupervised mask generation or segmentation.
Jointly Learning Several Tasks. Pre-train on large-scale and easily available text-region pairs datasets or detection using a unified detection and segmentation model.
Adopting Denoising Diffusion Models. Explore the feature representation of text-to-image diffusion models or utilize the generation ability to augment masks.
@@ 3.4 Open Vocabulary Segmentation
Although segmentation problems can be defined in various ways, such as semantic segmentation or panoptic segmentation, we categorize the methods based on their technical aspects.
@@
Utilizing VLMs to Leverage Recognition Capabilities. VLMs have shown remarkable performance in image classification by learning rich visual and linguistic representations. Therefore, it is natural to extend VLMs to semantic segmentation, which can be seen as a dense classification task. For instance, LSeg [14] aligns the text embeddings of category labels from a VLM language encoder with the dense embeddings of the input image. This enables LSeg to leverage the generalization ability of VLMs and segment objects that are not predefined but depend on the input texts. Following LSeg, several works propose methods to utilize the VLMs for open vocabulary segmentation tasks. Fusioner [252] utilizes self-attention operations to combine visual and language features in a transformer-based framework at an early stage. ZegFormer [114] decouples the problem into a class-agnostic segmentation task and a mask classification task. It uses the label embeddings from a VLM to classify the proposal masks and applies the CLIP-vision encoder to obtain language-aligned visual features for them. After that, more approaches are proposed to extract the rich knowledge in VLMs. SAN [253] attaches a lightweight side network to the pre-trained VLM to predict mask proposals and classification outputs. CAT-Seg [254] jointly aggregates the image and text embeddings of CLIP by fine-tuning the image encoder. Han et al. [255] develop an efficient framework that does not rely on the extra computational burden of the CLIP model. OPSNet [256] proposes several modulation modules to enhance the information exchange between the segmentation model and the VLMs. The modulation modules fuse the knowledge of the learned vision model and VLM, which improves the zero-shot performance in novel classes. MaskCLIP [48] inserts Relative Mask Attention (RMA) modules into a pre-trained CLIP model. It can utilize the CLIP features more efficiently. TagCLIP [257] proposes a trusty token module to explicitly predict pixels that contain objects (both base and novel) before classifying each pixel, which avoids the problem that models tend to misidentify pixels into novel categories. Recently, several works [258, 259, 260] directly fuse the frozen CLIP visual encoder. They fuse the learned CLIP score and prediction score to achieve better close-set and open-set recognition ability trade-offs. Same as the open vocabulary detection, there are still several improve space when adopting more advanced VLMs.
@@
Learning from Caption Data. Image caption data can provide weak supervision for identifying novel classes, as they expose potential novel category names. Like such approaches in open vocabulary object detection [11], image caption data is also explored in open vocabulary segmentation tasks also explore image caption data. OpenSeg [46] applies a region-word grounding loss to directly ground objects and nouns in the caption data. CGG [47] combines caption grounding and caption generation losses to fully exploit the knowledge in caption data. It leverages the role of object nouns in visual grounding and the mutual benefits of words in caption generation.
@@
Generating Pseudo Labels. Intuitively, providing the model with more data on novel categories can improve classification performance. MaskCLIP+ [120] modifies the CLIP-vision model by replacing the last pooling layer with a convolution layer, which produces dense feature maps. These feature maps are then used to generate pseudo labels for training a segmentation model. OVSeg [121] matches the proposed image regions with nouns in captions using CLIP to generate pseudo annotations. It also proposes a mask prompt tuning module to help CLIP adapt to masked images without changing their weights. XPM [122] follows a similar pseudo data generation procedure by aligning words in captions with regions in images to generate pseudo instance mask labels. However, these approaches are limited by the data annotations, since both mask and region caption are hard to collection. Maybe more automatic data generation [132, 42] can be used in the future.
@@
Training without Pixel-Level Annotations. Despite the absence of annotations on novel objects, most methods still require base mask annotations during training, which are costly and labor-intensive to obtain. To reduce the annotation burden, many recent works explore training segmentation models with merely weak supervision, such as image caption. GroupViT [125] proposes a semantic segmentation framework that leverages a grouping mechanism to automatically merge image patches with the same semantics. It trains the model with a contrastive loss on image-caption pairs and does not need pixel-level annotations. PACL [261] enhances the contrastive loss with a patch alignment objective that aligns the image patches and the CLS token of the captions. ViL-Seg [221] combines both contrastive loss and clustering loss. SegCLIP [126] further introduces a reconstruction loss and a superpixel-based KL loss. To compute the superpixel-based KL loss, it uses an unsupervised graph-based segmentation model [262] to generate pseudo labels. TCL [263] proposes a finer-grained contrastive loss, namely text-grounded contrastive loss, which explicitly aligns captions and regions. The model can segment the region that corresponds to a given text expression during inference. OVSegmentor [222] introduces masked entity completion and cross-image mask constituency tasks to improve the training efficiency. Mask-free OVIS [127] generates pseudo mask annotations with purely image-text pair data to assist the segmentation models. Since these methods are trained without mask supervision, the quality of segmentation results is low which make it hard to use in real application.
@@
Jointly Learning Several Tasks. Open vocabulary segmentation encompasses different tasks, including OVSS, OVIS, and OVPS. How to jointly learn multiple segmentation tasks in one model becomes a practical problem. X-Decoder [130] proposes a framework that can handle various tasks, including open vocabulary semantic segmentation, open vocabulary instance segmentation, and open vocabulary panoptic segmentation. It utilized a query-based segmentation architecture, pre-trains the model on a mixture of segmentation data and image-text pairs, and then fine-tuned or applied in zero-shot settings for downstream tasks. FreeSeg [131] also proposes a generic framework to tackle the three tasks in a unified manner. It designs an adaptive task prompt module and performs test time tuning on the learnable prompts to capture the task-specific features. POMP [264] first trains class prompts at a large vocabulary dataset (Imagenet-21K) and then transfers the prompts into multiple open vocabulary tasks. Moreover, if the model can learn from multiple tasks, it raises the question of whether it can benefit from multi-sourced data, e.g., detection and segmentation data. To handle this question, OpenSeeD [128] and OpenSD [129] jointly learn from segmentation and detection datasets. To bridge the task gap, they propose decoupled decoding frameworks, which decode foreground and background masks separately and generate masks for the bounding box proposals. One shortcoming of these approaches is extra computation costs that brought other tasks.
@@
Adopting Denoising Diffusion Models. Recently, diffusion-based generative models [265] have achieved remarkable success in text-based image generation suggests. There are mainly two ways to let diffusion models enhance open vocabulary tasks. First, The reality and diversity of the images generated by diffusion models suggest that the intermediate representations in the diffusion models may be highly aligned with natural language vocabularies. Inspired by this, ODISE [133] proposes a framework that leverages the middle representation of the diffusion model. The middle representations contain rich semantic information for both base and novel classes. Thus, ODISE can perform open vocabulary segmentation by training a decoder head on the representations. [266] follows ODISE to use the intermediate representations of diffusion models. Another way to incorporate diffusion models into the open vocabulary setting is to take advantage of the image and mask generation ability [139, 132, 267]. OVDiff [132] proposes a prototype-based method to tackle open vocabulary semantic segmentation. It uses the diffusion model to generate images for various categories and treat them as prototypes. The method does not need training. During testing, the input images are compared with these generated prototypes, and the best-matching one is the predicted class. Meanwhile, several works [139, 267] use the diffusion model to generate images and masks for the rare classes to augment data. Thus, the segmenter can be trained in a more balanced manner or even totally using generated data. Despite the inference time of these methods is quite limited, it still has room to explore joint generation and segmentation in one framework.